OpenStudy (anonymous):

Given that f(x) = 5x - 10 and g(x) = x + 3, solve for f(g(x)) when x = -1

OpenStudy (anonymous):

You want f(g(1)). First compute g(1). Then whatever you get for g(1), you will use that number to find f(that number).

OpenStudy (anonymous):

so g(-1)=-1+3 g=-2?

OpenStudy (anonymous):

Sorry, you wanted at -1. so compute f(g(-1)).

OpenStudy (anonymous):

g(-1) = 2 so f(2) = ??

OpenStudy (anonymous):

f(2) = 5(2) - 10 = 0.

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

I get it... Thankyou!